If you want to mix 50lb of nuts worth $2.00 per lb with some nuts worth $6.00 per lb to make a mixture worth $5.00 per lb. How many pounds of $6.00 nuts must you use?

That is the question. How do I turn this into an equation or whatever I need to do?

Answer by ankor@dixie-net.com(22740)

If you want to mix 50lb of nuts worth $2.00 per lb with some nuts worth $6.00 per lb to make a mixture worth $5.00 per lb.
How many pounds of $6.00 nuts must you use?
:
A typical mixture equation, learn this method and you can do most of these in the future
:
Let x = no. of lb of $6 nut required
then you know the resulting mixture = (x + 50)
:
2(50) + 6x = 5(x + 50)
100 + 6x = 5x + 250
6x - 5x = 250 - 100
x = 150 lb of $6 nuts
:
:
Check your solution in the original equation
2(50) + 6(150) = 5(150 + 50)
100 + 900 = 5(200); confirms the solution
:
:
That's not so hard, is it?